What volume of a 0.75 mol/L solution of aluminum iodide contains in 4.5 x 10-2 moles?
1
Expert's answer
2020-10-21T13:21:17-0400

Cm=n/V, V= n/Cm=4.5 x 10-2/0.75 =0.06 L.